OpenStudy (anonymous):

Show work 8c+5=4c-11

OpenStudy (anonymous):

8c - 4c = -11 - 5 4c = -16 c = -4

OpenStudy (saifoo.khan):

8c - 4c = -11 -5 4c = -16 c = -16/4 c = -4